OpenStudy (anonymous):

given that SU=x+1 and RT=2x-3, find the value of x in the isosceles trapezoid

OpenStudy (mathmale):

Isn't there a diagram with this problem? If so, won't you please post it?

OpenStudy (mathmale):

The key word/idea here is that the trapezoid is isosceles. What does that mean to you? To me it means that side SU = side RT. Thus, side SU = x+1 = 2x-3 = side RT. What next? What's your goal here?

OpenStudy (anonymous):

to find x?

OpenStudy (mathmale):

Yes. And what does "isosceles" mean when we're discussing trianges? trapezoids?

OpenStudy (anonymous):

two sides are equal?...im not good with this stuff

OpenStudy (mathmale):

Yes, an isosceles triangle has two equal sides and one more side that may or may not be equal to one of the first two sides. So, an isosceles trapezoid .... what?

OpenStudy (anonymous):

ummm....I don't know :/

OpenStudy (anonymous):

ive understood everything but this part so far

OpenStudy (mathmale):

Refer to the diagram that came with this problem. Since the trapez. is "isosceles," two of its sides are equal. Which ones?

OpenStudy (anonymous):

RU and ST?

OpenStudy (mathmale):

Yes. Now, knowing that RU=ST, can you say anything about the relationship between SU and RT? This is the key to solving this particular problem.

OpenStudy (mathmale):

JN: It's been a long while since I last went through geometric proofs. However, we're given quite a lot of info about this isosceles trapezoid. Perhaps someone more up to date on such proofs could go into more detail than I. However, in the end, you'll need to show that SU=RT, and thus, side SU = x+1 = 2x-3 = side RT; solve x+1 = 2x - 3 for x and you'll be done.

OpenStudy (anonymous):

ok thanks

OpenStudy (mathmale):

Given x+1 = 2x - 3, subtract x from both sides. Then add 3 to both sides. x=??

OpenStudy (anonymous):

4

OpenStudy (anonymous):

i think i did something wrong there

OpenStudy (anonymous):

but anyways i got it i accidently put the wrong number thanks